# Generalizable Sleep Staging via Multi-level Domain Alignment

### 13 Dec 2023 v1

@@ Introduction
Sleep plays an important role in human health. Sleep staging refers to the classification of sleep into different sleep stages, which is crucial to identifying sleep problems and disorders in humans. Clinically, sleep stages are scored by doctors or experts using electrical activity recorded from sensors attached to different parts of the body. A set of signals from these sensors is called a polysomnogram (PSG), consisting of multiple physiological signals recorded, such as electroencephalogram (EEG) and electrooculogram (EOG). According to the American Academy of Sleep Medicine (AASM) sleep standard (Iber et al. 2007), PSG is usually segmented into 30-second epochs, which are manually classified into five different sleep stages (Wake, N1, N2, N3, and REM) by experts. In recent years, with the development of data-driven deep learning techniques, many deep learning models (Supratak et al. 2017; Seo et al. 2020; Perslev et al. 2019; Phan et al. 2021; Jia et al. 2020, 2021) have been proposed to solve the task of automatic sleep staging. They are implemented with different network structures based on sequence-to-sequence framework, obtaining good performance in sleep staging from PSG recordings. However, most of existing methods adopt the intra-dataset scheme, where the training data and testing data are from the same dataset, ignoring the discrepancies between various datasets and making it difficult to be generalized to unseen sleep staging datasets well (Anido-Alonso and Alvarez-Estevez 2022). The discrepancies among multiple datasets could be caused by many factors, such as heterogeneous patient population, different signal channel, different data collection equipment types or manners, or different medical environments. In clinic, a sleep staging model should better be generalized to unseen datasets of any new populations in any environment. Some methods (Fan et al. 2022; Yoo, Lee, and Kang 2021) illustrate that the training and testing samples are not drawn from the same probability distribution which cause performance deterioration when directly applying models trained on source datasets to target datasets. They adopt the ideas of Domain Adaptation (DA) to solve it, but the generalization problem still exist because DA methods assume that the target sleep staging database is accessible. In order to solve the generalizing-to-unseen-database problem of automatic sleep staging, we introduce the idea of domain generalization (DG) into the sleep staging task: A domain is composed of data that is sampled from a join distribution of input space and label space. The goal of domain generalization is to learn a model from one or several different but related domains (i.e., diverse training datasets) that will perform well on unseen testing domains (Wang et al. 2023a). The unseen domain is inaccessible in training procedure. Inspired by some studies (Koh et al. 2021; Robey, Pappas, and Hassani 2021) exploring hospital-level DG in medical imaging, we set a single sleep staging dataset (usually from one hospital) as a domain and set the discrepancies between different datasets as domain shift. Then multiple datasets used for training can be regarded as source domains and the dataset used for testing can be seen as target domain or unseen domain. Our goal is to learn a model from several source domains (several sleep datasets) that can perform well at the unseen domain (unseen sleep dataset). We name this task as generalizable sleep staging. There have been some existing DG studies (Muandet, Balduzzi, and Schölkopf 2013; Li et al. 2018a, b; Matsuura and Harada 2020; Zhou et al. 2020) in other applications, which mainly focus on learning domain-invariant feature representations by sample-level feature distribution alignment to solve the sample-level classification problem, referring to mapping single sample into the corresponding label. However, different from sample-level classification problem, sleep staging is a sequence-to-sequence classification problem, which maps a sequence of samples into the corresponding sequence of labels. According to AASM standard (Iber et al. 2007), not only local salient waveforms within each epoch but also transition patterns of sleep stages between neighbor epochs play critical roles in sleep staging. Meantime, there have been some studies (Seo et al. 2020; Eldele et al. 2021; Phan et al. 2021) proving the importance of both local intra-epoch features and global inter-epoch features for sleep staging. Therefore, it is critical to combine intra-epoch feature (epoch-level) alignment and inter-epoch feature (sequence-level) alignment in alignment process, to solve the generalization problem of sleep staging. In this paper, we introduce DG to sleep staging and propose a framework, SleepDG, to solve the generalization problem of sleep staging. Our contributions are as follows:
@@
DG (Gulrajani and Lopez-Paz 2021; Zhou et al. 2023; Wang et al. 2023a) considers the scenarios where there are domain shift among different domains, which is different from domain adaptation (DA) (Wang and Deng 2018; Zhao, Yan, and Lu 2021) because target domain in DG is inaccessible. The DG problem was formally introduced by Blanchard, Lee, and Scott (2011); Muandet, Balduzzi, and Schölkopf (2013) as a machine learning problem. Learning domain-invariant feature representations by minimizing the feature distribution divergence is a common and effective method for domain generalization. Some studies minimized the feature distribution divergence by minimizing the maximum mean discrepancy (MMD) (Tzeng et al. 2014; Wang et al. 2018, 2020), second order correlation (Sun, Feng, and Saenko 2016; Sun and Saenko 2016; Peng and Saenko 2018), Wasserstein distance (Zhou et al. 2020) and Similarity Metric (Dou et al. 2019) in deep neural network. Some studies minimized the feature distribution divergence by adopting Domain-adversarial neural network (DANN) (Li et al. 2018a) for learning domain-invariant features, which was originally proposed by Ganin and Lempitsky (2015); Ganin et al. (2016) for DA. Most of above methods learn domain-invariant features by minimizing single-sample features divergence, neglecting the crucial role of inter-epoch features in sleep staging, which needs to be improved.
@@
The feature alignment refers to mapping source data from different datasets into one shared feature space in which $\mathrm{P}(H^{S_{i}},Y^{S_{i}})=\mathrm{P}(H^{S_{j}},Y^{S_{j}})$ if $i\neq j$. We can decompose joint distribution $\mathrm{P}(H^{S_{i}},Y^{S_{i}})$ into a marginal distribution and a conditional distribution by $\mathrm{P}(H^{S_{i}},Y^{S_{i}})=\mathrm{P}(H^{S_{i}})\mathrm{P}(Y^{S_{i}}|H^{S_{i}})$. According to AASM standard (Iber et al. 2007), sleep experts adopt a uniform rules for manual sleep staging, so we assume that the conditional distributions are the same in different sleep staging datasets. In other words, we make a prior assumption that $\mathrm{P}(Y^{S_{i}}|H^{S_{i}})=\mathrm{P}(Y^{S_{j}}|H^{S_{j}})$, if $i\neq j$. Under such a prior assumption, we only need to align the marginal distribution $\mathrm{P}(H^{S_{i}})$. Sleep staging is a sequence-to-sequence classification and both local intra-epoch features and global context features among sequential epochs are important for sleep staging (Seo et al. 2020; Eldele et al. 2021; Phan et al. 2021). Inspired by this, we design a Multi-level Feature Alignment method consisting of Epoch-level Feature Alignment and Sequence-level Feature Alignment. The alignment of sequence features $\mathrm{\mathbf{h}}^{S_{i}}_{j}=\{h^{S_{i}}_{j,1},h^{S_{i}}_{j,2},...,h^{S_{i}}_{j,L}\}$ could be divided to reducing marginal distribution on single epoch features and conditional distribution between epochs in a sleep sequence:
@@
Some existing methods minimize the feature distribution divergence by minimizing the maximum mean discrepancy (MMD) (Tzeng et al. 2014; Pan et al. 2010; Wang et al. 2018, 2020) and other methods minimize the domain discrepancy by minimizing the second order correlation (Sun, Feng, and Saenko 2016; Sun and Saenko 2016; Peng and Saenko 2018). Inspired by above, we utilize both the first-order statistics (expectation) and the second-order statistics (covariances) as the measure of distribution to minimize the domain discrepancy. We set $F^{i}$ as the set of all features in the source domain $S_{i}$, so any $h^{S_{i}}_{j,k}\in F^{i}$. The first-order statistics discrepancy is computed through the following equation:
@@ V. P2018 200 995 150 C3-M2 E1-M2
We compared SleepDG with BASE and other classical DG methods in sleep staging: BASE can be regarded as SleepDG which minimizes $\mathcal{L}_{\mathrm{classify}}$ but without any reconstruction and DG loss. MMD (Maximum Mean Discrepancy) (Tzeng et al. 2014; Li et al. 2018a) is a DG method matching the MMD of feature distributions. CORAL (Sun and Saenko 2016) matches the covariance of feature distributions. IRM (Invariant Risk Minimization) (Arjovsky et al. 2019) enforces the optimal classifier to be the same across all domains. DANN (Domain-Adversarial Neural Networks) (Ganin and Lempitsky 2015; Ganin et al. 2016) employs an adversarial network to match feature distributions. We implemented these DG methods based on BASE and DeepDG toolkit (Wang and Lu 2021), and the performance comparison is shown in Tab. 3. SleepDG achieves the best performance compared with other DG methods. BASE performs the worst (68.76% in average ACC and 63.34% in average MF1), indicating that a model without DG loss cannot minimize the domain shift among different domains. MMD performs better compared with BASE, about 3.0% higher in average ACC and 2.7% higher in average MF1, suggesting minimizing maximum mean discrepancy can extract domain-invariant features to a degree. CORAL performs the best in existing DG methods (72.27% in average ACC and 67.52% in average MF1), which indicates that feature alignment based on second-order statistics (covariance) can achieve good cross-domain generalization. IRM performs better than MMD but worse than CORAL (72.09% in average ACC and 66.95% in average), indicating that invariant risk minimization has no obvious advantage over feature alignment methods. The domain-adversarial model of DANN achieves the least improvement compared with BASE (71.12% v.s. 68.76% in average ACC). It matches the finding in Wang et al. (2023a): although domain adversarial training often achieves better performance in DA, it is difficult to make a significant improvement in DG. Compared with these classical DG methods, SleepDG can learn domain-invariant features better by combining Epoch-level Feature Alignment and Sequence-level Feature Alignment.
@@ Acknowledgments
This work was supported by the Key Program of the Natural Science Foundattion of Zhejiang Province, China (No. Z24F020009), STI 2030 Major Projects (No. 2021ZD0200400) and Natural Science Foundation of China (No. 61925603). The corresponding author is Dr. Sha Zhao.

### 12 Jan 2024 v2

@@ Introduction
Sleep plays an important role in human health. Sleep staging refers to the classification of sleep into different sleep stages, which is crucial to identifying sleep problems and other disorders in humans (He et al. 2018). Clinically, sleep stages are scored by doctors or experts using electrical activity recorded from sensors attached to different parts of the body. A set of signals from these sensors is called a polysomnogram (PSG), consisting of multiple physiological signals recorded, such as electroencephalogram (EEG) and electrooculogram (EOG). According to the American Academy of Sleep Medicine (AASM) sleep standard (Iber et al. 2007), PSG is usually segmented into 30-second epochs, which are manually classified into five different sleep stages (Wake, N1, N2, N3, and REM) by experts. In recent years, with the development of deep learning techniques (Sun et al. 2016; Pan et al. 2018; Yi et al. 2021; Wang et al. 2023c), many deep learning models (Phan and Mikkelsen 2022) have been proposed to solve the task of automatic sleep staging. They are implemented with different network structures based on sequence-to-sequence framework, obtaining good performance in sleep staging from PSG recordings. However, most of existing methods adopt the intra-dataset scheme, where the training data and testing data are from the same dataset, ignoring the discrepancies between various datasets and making it difficult to be generalized to unseen sleep staging datasets well (Anido-Alonso and Alvarez-Estevez 2022). The discrepancies among multiple datasets could be caused by many factors, such as heterogeneous patient population, different signal channel, different data collection equipment types or manners, or different medical environments. In clinic, a sleep staging model should better be generalized to unseen datasets of any new populations in any environment. Some methods (Yoo, Lee, and Kang 2021; Fan et al. 2022; Wang et al. 2022) illustrate that the training and testing samples are not drawn from the same probability distribution which cause performance deterioration when directly applying models trained on source datasets to target datasets. They adopt the ideas of Domain Adaptation (DA) to solve it, but the generalization problem still exist because DA methods assume that the target sleep staging database is accessible. In order to solve the generalizing-to-unseen-database problem of automatic sleep staging, we introduce the idea of domain generalization (DG) into the sleep staging task: A domain is composed of data that is sampled from a join distribution of input space and label space. The goal of domain generalization is to learn a model from one or several different but related domains (i.e., diverse training datasets) that will perform well on unseen testing domains (Wang et al. 2023a). The unseen domain is inaccessible in training procedure. Inspired by some studies (Koh et al. 2021; Robey, Pappas, and Hassani 2021) exploring hospital-level DG in medical imaging, we set a single sleep staging dataset (usually from one hospital) as a domain and set the discrepancies between different datasets as domain shift. Then multiple datasets used for training can be regarded as source domains and the dataset used for testing can be seen as target domain or unseen domain. Our goal is to learn a model from several source domains (several sleep datasets) that can perform well at the unseen domain (unseen sleep dataset). We name this task as generalizable sleep staging. There have been some existing DG studies (Muandet, Balduzzi, and Schölkopf 2013; Li et al. 2018a, b; Matsuura and Harada 2020; Zhou et al. 2020) in other applications, which mainly focus on learning domain-invariant feature representations by sample-level feature distribution alignment to solve the sample-level classification problem, referring to mapping single sample into the corresponding label. However, different from sample-level classification problem, sleep staging is a sequence-to-sequence classification problem, which maps a sequence of samples into the corresponding sequence of labels. According to AASM standard (Iber et al. 2007), not only local salient waveforms within each epoch but also transition patterns of sleep stages between neighbor epochs play critical roles in sleep staging. Meantime, there have been some studies (Eldele et al. 2021; Phan et al. 2021) proving the importance of both local intra-epoch features and global inter-epoch features for sleep staging. Therefore, it is critical to combine intra-epoch feature (epoch-level) alignment and inter-epoch feature (sequence-level) alignment in alignment process, to solve the generalization problem of sleep staging. In this paper, we introduce DG to sleep staging and propose a framework, SleepDG, to solve the generalization problem of sleep staging. Our contributions are as follows:
@@
DG (Gulrajani and Lopez-Paz 2021; Zhou et al. 2023; Wang et al. 2023a) considers the scenarios where there are domain shift among different domains, which is different from domain adaptation (DA) (Wang and Deng 2018; Zhao, Yan, and Lu 2021) because target domain in DG is inaccessible. The DG problem was formally introduced by Blanchard, Lee, and Scott (2011); Muandet, Balduzzi, and Schölkopf (2013) as a machine learning problem. Learning domain-invariant feature representations by minimizing the feature distribution divergence is a common and effective method for domain generalization. Some studies minimized the feature distribution divergence by minimizing the maximum mean discrepancy (MMD) (Tzeng et al. 2014; Wang et al. 2018, 2020), second order correlation (Sun, Feng, and Saenko 2016; Sun and Saenko 2016), Wasserstein distance (Zhou et al. 2020) and Similarity Metric (Dou et al. 2019) in deep neural network. Some studies minimized the feature distribution divergence by adopting Domain-adversarial neural network (DANN) (Li et al. 2018a) for learning domain-invariant features, which was originally proposed by Ganin and Lempitsky (2015) for DA. Most of above methods learn domain-invariant features by minimizing single-sample features divergence, neglecting the crucial role of inter-epoch features in sleep staging, which needs to be improved.
@@
The feature alignment refers to mapping source data from different datasets into one shared feature space in which $\mathrm{P}(H^{S_{i}},Y^{S_{i}})=\mathrm{P}(H^{S_{j}},Y^{S_{j}})$ if $i\neq j$. We can decompose joint distribution $\mathrm{P}(H^{S_{i}},Y^{S_{i}})$ into a marginal distribution and a conditional distribution by $\mathrm{P}(H^{S_{i}},Y^{S_{i}})=\mathrm{P}(H^{S_{i}})\mathrm{P}(Y^{S_{i}}|H^{S_{i}})$. According to AASM standard (Iber et al. 2007), sleep experts adopt a uniform rules for manual sleep staging, so we assume that the conditional distributions are the same in different sleep staging datasets. In other words, we make a prior assumption that $\mathrm{P}(Y^{S_{i}}|H^{S_{i}})=\mathrm{P}(Y^{S_{j}}|H^{S_{j}})$, if $i\neq j$. Under such a prior assumption, we only need to align the marginal distribution $\mathrm{P}(H^{S_{i}})$. Sleep staging is a sequence-to-sequence classification and both local intra-epoch features and global context features among sequential epochs are important for sleep staging (Eldele et al. 2021; Phan et al. 2021). Inspired by this, we design a Multi-level Feature Alignment method consisting of Epoch-level Feature Alignment and Sequence-level Feature Alignment. The alignment of sequence features $\mathrm{\mathbf{h}}^{S_{i}}_{j}=\{h^{S_{i}}_{j,1},h^{S_{i}}_{j,2},...,h^{S_{i}}_{j,L}\}$ could be divided to reducing marginal distribution on single epoch features and conditional distribution between epochs in a sleep sequence:
@@
Some existing methods minimize the feature distribution divergence by minimizing the maximum mean discrepancy (MMD) (Tzeng et al. 2014; Wang et al. 2018, 2020) and other methods minimize the domain discrepancy by minimizing the second order correlation (Sun, Feng, and Saenko 2016; Sun and Saenko 2016). Inspired by above, we utilize both the first-order statistics (expectation) and the second-order statistics (covariances) as the measure of distribution to minimize the domain discrepancy. We set $F^{i}$ as the set of all features in the source domain $S_{i}$, so any $h^{S_{i}}_{j,k}\in F^{i}$. The first-order statistics discrepancy is computed through the following equation:
@@ V. P2018 200 995 150 C3-M2 E1-M2
We compared SleepDG with BASE and other classical DG methods in sleep staging: BASE can be regarded as SleepDG which minimizes $\mathcal{L}_{\mathrm{classify}}$ but without any reconstruction and DG loss. MMD (Maximum Mean Discrepancy) (Tzeng et al. 2014; Li et al. 2018a) is a DG method matching the MMD of feature distributions. CORAL (Sun and Saenko 2016) matches the covariance of feature distributions. IRM (Invariant Risk Minimization) (Arjovsky et al. 2019) enforces the optimal classifier to be the same across all domains. DANN (Domain-Adversarial Neural Networks) (Ganin and Lempitsky 2015) employs an adversarial network to match feature distributions. We implemented these DG methods based on BASE and DeepDG toolkit (Wang et al. 2023a), and the performance comparison is shown in Tab. 3. SleepDG achieves the best performance compared with other DG methods. BASE performs the worst (68.76% in average ACC and 63.34% in average MF1), indicating that a model without DG loss cannot minimize the domain shift among different domains. MMD performs better compared with BASE, about 3.0% higher in average ACC and 2.7% higher in average MF1, suggesting minimizing maximum mean discrepancy can extract domain-invariant features to a degree. CORAL performs the best in existing DG methods (72.27% in average ACC and 67.52% in average MF1), which indicates that feature alignment based on second-order statistics (covariance) can achieve good cross-domain generalization. IRM performs better than MMD but worse than CORAL (72.09% in average ACC and 66.95% in average), indicating that invariant risk minimization has no obvious advantage over feature alignment methods. The domain-adversarial model of DANN achieves the least improvement compared with BASE (71.12% v.s. 68.76% in average ACC). It matches the finding in Wang et al. (2023a): although domain adversarial training often achieves better performance in DA, it is difficult to make a significant improvement in DG. Compared with these classical DG methods, SleepDG can learn domain-invariant features better by combining Epoch-level Feature Alignment and Sequence-level Feature Alignment.
@@ Acknowledgments
This work was supported by the Key Program of the Natural Science Foundattion of Zhejiang Province, China (No. Z24F020009), STI 2030 Major Projects (No. 2021ZD0200400) and Natural Science Foundation of China (No. 61925603). The corresponding authors are Dr. Sha Zhao and Dr. Gang Pan.